OpenStudy (anonymous):

PLEASE HELP!! simplifying 13 ---- 56

OpenStudy (anonymous):

I dont think you can

OpenStudy (anonymous):

@WhiteFlower there is nothing you can do here

OpenStudy (anonymous):

Good job you solved it!!!

OpenStudy (anonymous):

\[\frac{13}{56}=\frac{13}{2^3\times 7}\] there are no common factors, so you cannot reduce in any way

ganeshie8 (ganeshie8):

is it \(1\frac{3}{56}\) ?